OpenStudy (anonymous):

Linda is studying the sale of a particular brand of cereals from the year 1993 to 2004. She writes the following function to model the sale of the cereal S(t), in million dollars, after t years: S(t) = t2 + 5t + 52 Part A: What does the y-intercept of the graph of the function represent? (4 points) Part B: What is the reasonable domain of the graph of the function? (3 points) Part C: What is the average rate of change of the sale of the cereal from the first year to the fourth year? Show your work. (3 points)

OpenStudy (anonymous):

ok so for this one i graphed it and i see that the y intercept is 0,52 but im not sure what it represents

OpenStudy (anonymous):

ill post what the graph looks like

OpenStudy (anonymous):

what do you think pawtpie

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

The y intercept would have to be the value when t = 0

OpenStudy (anonymous):

i understand that but i have no idea what it would represent maybe the starting price

OpenStudy (anonymous):

@Pawtpie so when t=0 should be how much cereal she starts with I would think

OpenStudy (anonymous):

i mean the starting amount of cereal

OpenStudy (anonymous):

I think that would be the starting amount of 52 in millions

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

and for the domain i know that the domain is all real numbers am i right

OpenStudy (anonymous):

@melancholymushroom yes

OpenStudy (anonymous):

ok so i would have to come up with something like this x<?<y right

OpenStudy (anonymous):

I think so yes

OpenStudy (anonymous):

ok so how exactly do i do this do i take all the x values

OpenStudy (anonymous):

Maybe the x value increases by 1 each year so it would be the difference in years.

OpenStudy (anonymous):

so 1 less than or equal x ? somthing like that

OpenStudy (anonymous):

Well if we're looking at the years between 1993 and 2004 that's 11 years

OpenStudy (anonymous):

so maybe it's 0<x<11

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hold on small typo actually

OpenStudy (anonymous):

the problem is that i need it to be more than or equal too but how do i write that out on my keyboard

OpenStudy (anonymous):

actually might be 0<=x<=11

OpenStudy (anonymous):

I think we would just write the domain as {0, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11} instead

OpenStudy (anonymous):

ok and for part c i do average rate of change so that would be is it years/amount of change

OpenStudy (anonymous):

x-y/x-y

OpenStudy (anonymous):

is the bottom supposed to be the years or the difference in numbers

OpenStudy (anonymous):

Maybe we just use the formula for the slope between two points is what you're saying

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

I guess we just get S(1) and S(4) for t to get the 2 ordered pairs and then get the slope

OpenStudy (anonymous):

for the average rate of change i know its supposed to be the difference in the amount of year so 4-1 and the difference in teh amount of cereal sold which was 88-?

OpenStudy (anonymous):

im trying to find the second point

OpenStudy (anonymous):

All I did to get the points was find S(1) and S(4)

OpenStudy (anonymous):

ok then i guess im doing it wrong what do i do to find the average rate of change

OpenStudy (anonymous):

cuz i have to find the average rate of change from year 4 to 1

OpenStudy (anonymous):

Ok first we need 2 ordered pairs to find the slope from them

OpenStudy (anonymous):

What I did was I got S(1) and S(4)

OpenStudy (anonymous):

ok two points were 1,58 and 4,88

OpenStudy (anonymous):

I got (4, 78) for the second one

OpenStudy (anonymous):

oh whoops i meant 78 typo

OpenStudy (anonymous):

Ok so it seems that all we have to do is find the distance between the two points: \[Slope = \frac{ rise }{ run }=\frac{ y2 - y1 }{ x2 - x1 }\]

OpenStudy (anonymous):

For (1, 58) and (4, 78)

OpenStudy (anonymous):

so would it be 1/10 or 10/1

OpenStudy (anonymous):

I'll check real quick

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Seems I got 20/3

OpenStudy (anonymous):

\[\frac{ 78 - 58 }{ 4 - 1 } = \frac{ 20 }{ 3 }\]

OpenStudy (anonymous):

oh whoops i put 88 instead of 78 i see now

OpenStudy (anonymous):

alright thats pretty much it thanks for all the help i really appreciate it

OpenStudy (anonymous):

No problemo :)

OpenStudy (anonymous):

btw nice profile pic lol what is it

OpenStudy (anonymous):

Honestly have no idea, I just choose from the Google images of "derp".

OpenStudy (anonymous):

lol good to know thanks for all the help

OpenStudy (anonymous):

:3

OpenStudy (anonymous):

No problem again lol